OpenStudy (anonymous):

What is the x-coordinate of the vertex of this parabola? y = -3x^2+12x-5

OpenStudy (anonymous):

y = -3x² +12x - 5 y = -3(x² + 4x) - 5 y = -3(x² + 4x + 4 - 4) - 5 y = -3(x² + 4x + 4) + 12 - 5 y = -3(x + 2)² + 7 x-coordinate is (-2 , 7)

OpenStudy (april115):

what is a parabola?

OpenStudy (anonymous):

Answer Choices... A.)-3 B.)2 C.)12

OpenStudy (april115):

is it B)

OpenStudy (anonymous):

Thats what im wondering..

OpenStudy (april115):

me too....

OpenStudy (april115):

well, let's see if @HairtUB knows the answer.

OpenStudy (anonymous):

^Right

OpenStudy (april115):

ok=]

OpenStudy (anonymous):

Ok wait, yeah it's B. I made a mistake in my calculations. Here's the correct version: y = -3x² +12x - 5 y = -3(x² - 4x) - 5 y = -3(x² - 4x + 4 - 4) - 5 y = -3(x² - 4x + 4) + 12 - 5 y = -3(x - 2)² + 7 coordinate is (-2 , 7) x-coordinate is 2.

OpenStudy (anonymous):

^Thanks

OpenStudy (april115):

ok, i was right, thank u

OpenStudy (anonymous):

you were :D

OpenStudy (april115):

:D

OpenStudy (anonymous):

No prob :) And sorry for the confusion :/

OpenStudy (april115):

its okay:)